OpenStudy (anonymous):

Given the functions a(x) = 3x − 12 and b(x) = x − 9, solve a[b(x)] and select the correct answer below. a[b(x)] = 3x2 − 21 a[b(x)] = 3x2 − 39 a[b(x)] = 3x − 21 a[b(x)] = 3x − 39

OpenStudy (anonymous):

anyy help

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

@Vocaloid i am your 200th fan the least u could do is help me out

Vocaloid (vocaloid):

a(x) = 3x - 12 a((b(x)) = 3*b(x) - 12 = 3(x-9) - 12

OpenStudy (anonymous):

what do i do

Vocaloid (vocaloid):

a(x) = 3x - 12, right? a(b(x)) means replace "x" with b(x), so a(b(x)) = 3*b(x) - 12 now, we know that b(x) = x - 9, so we can replace "b(x)" with (x-9), giving us a(b(x)) = 3(x-9)-12 distribute the 3 and simplify

OpenStudy (anonymous):

3x-39 ?? @Vocaloid

Vocaloid (vocaloid):

yes